OpenStudy (anonymous):

The area of a sector of a circle with a radius of 4 centimeters is 2.512 square centimeters. The estimated value of π is 3.14. The measure of the angle defining the sector is °

OpenStudy (campbell_st):

the formula for area of a sector is \[A = \frac{x}{360} \times \pi r^2\] so you can rewrite the formula as \[\frac{A \times 360}{\pi \times r^2} = x\] substitute your values to find x

OpenStudy (anonymous):

would it be 18?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

you are welcome fill free to fan